At the end of 2020, Payne Industries had a deferred tax asset account with a balance of $105 million attributable to a temporary book-tax difference of $420 million in a liability for estimated expenses. At the end of 2021, the temporary difference is $320 million. Payne has no other temporary differences and no valuation allowance for the deferred tax asset. Taxable income for 2021 is $756 million and the tax rate is 25%.

Required:
1. Prepare the journal entry(s) to record Payne’s income taxes for 2021, assuming it is more likely than not that the deferred tax asset will be realized.
2. Prepare the journal entry(s) to record Payne’s income taxes for 2021, assuming it is more likely than not that only one-fourth of the deferred tax asset ultimately will be realized.
Answer #1

1. Deferred tax asset at the beginning of year in 2021 = $105 million

Deferred tax asset at the end of year 2021 = $320 * 25% = $80 million

Deferred tax asset for the year = $80 million - $105 million = - $25 million

Tax Expense = Tax payable + Deferred tax asset = ($756 * 25%) + $25 million = $214 million

Journal Entry(Amount in millions)

Tax Expense Dr $214

To Deferred Tax Asset $25

To Tax Payable $189

(Being Payne's income taxes recorded for 2021)

2. 1/4th of Deferred tax asset in 2021 = 1/4 of $80 million = $20 million

Valuation Allowance = $80 million - $20 million = $60 million

Journal Entry (Amount in millions)

Tax Expense Dr $214

To Deferred Tax Asset $25

To Tax Payable $189

(Being Payne's income taxes recorded for 2021)

Income tax expense Dr $60

To Valuation Allowance - Deferred tax asset $60

(Being valuation allowance recorded as tax Expense)